Finney Displays designs and manufactures displays used in mobile devices. Serious flooding throughout the region affected Finney Displays​' facilities. Inventory was completely​ ruined, and the​ company's computer​ system, including all accounting​ records, was destroyed.

Before the disaster recovery specialists clean the​ buildings, Heather Bailey​, the company ​controller, is anxious to salvage whatever records she can to support an insurance claim for the destroyed inventory. She is standing in what is left of the Accounting Department wtih Tad Myers​, the cost accountant. ​"I didn't know mud could smell so​ bad," Tad says.​ "What should I be looking​ for?" ​"Don't worry about beginning inventory​ numbers," responds Heather. "We'll get them from last​year's annual report. We need​ first-quarter cost​ data." "I was working on the​ first-quarter results just before the storm​ hit," Tad says.​ "Look, my​report's still in my desk drawer. But all I can make out is that for the first​ quarter, material purchases were $524,000 and that direct​ labor, manufacturing overhead​ (other than indirect​materials), and total manufacturing costs to account for were $545,000​; $218,000​; and$1,508,000​, respectively.​ Wait, and cost of goods available for sale was $1,615,000​." ​"Great," says Heather. ​"I remember that sales for the period were approximately $1.6million. Given our gross profit of 15​%, ​that's all you should​ need." Tad is not sure about​ that, but decides to see what he can do with this information.

Raw​ materials, $85,000

Work in​ process, $187,000

Finished​ goods, $209,000

He remembers several schedules he learned in college that may help him get started.

Start by determining the ending inventory of raw materials by calculating the direct materials used.

Finney Displays

Calculation of Direct Materials Used

For Current Year

Plus:

Less:

​Next, determine the ending inventory of work in process by calculating the cost of goods manufactured.

Finney Displays

Calculation of Cost of Goods Manufactured

For Current Year

Plus:

Less:

​Finally, determine the ending inventory of finished goods by calculating the cost of goods manufactured.

Finney Displays

Calculation of Cost of Goods Sold

For Current Year

Plus:

Less:
Answer #1

Part 1

Finney Displays

Calculation of Direct Materials Used

For Current Year

Beginning raw materials inventory

85000

Plus: Purchases of direct materials

524000

Materials available for use

609000

Less: Ending raw materials inventory

51000

Direct materials used

558000

Part 2

Finney Displays

Calculation of Cost of Goods Manufactured

For Current Year

Beginning work in process inventory

187000

Plus: Manufacturing costs incurred

Direct labor

545000

Direct materials used

558000

Manufacturing overhead

218000

Total manufacturing costs to account for

1508000

Less: Ending work in process inventory

102000

Cost of goods manufactured

1406000

Part 3

Finney Displays

Calculation of Cost of Goods Sold

For Current Year

Beginning finished goods inventory

209000

Plus: Cost of goods manufactured

1406000

Cost of goods available for sale

1615000

Less: Ending finished goods inventory

255000

Cost of goods sold

$1360000

Cost of goods sold = sales * (1-gross profit percentage) = 1600000*(1-15%) = $1360000

Ending finished goods inventory = Cost of goods available for sale - Cost of goods sold = 1615000-1360000 = $255000

Cost of goods manufactured = Cost of goods available for sale - Beginning finished goods inventory = 1615000-209000 = 1406000

Ending work in process = Total manufacturing costs to account for - Cost of goods manufactured = 1508000-1406000 = 102000

Total manufacturing cost incurred = Total manufacturing costs to account for – beginning WIP = 1508000 -187000 = 1321000

Direct materials used = Total manufacturing cost incurred- direct labor – manufacturing overhead = 1321000-545000-218000 = 558000

Ending raw materials inventory = beginning direct materials + purchases – direct materials used = 85000+524000-558000 = 51000